# Robust Inference for Non-Linear Regression Models with Applications in Enzyme Kinetics

### v1

@@ 4.1 Asymptotic properties
In order to find the asymptotic distributions of the MDPDEs of the parameters of the MM model, we follow the notation of Section 2.2 to get $\bm{\mu(\beta)}=\left(\frac{\beta_{1}x_{1}}{\beta_{2}+x_{1}},\ldots,\frac{\beta_{1}x_{n}}{\beta_{2}+x_{n}}\right)^{T}$ and hence
$\bm{\dot{\mu}(\beta)}=\begin{bmatrix}\frac{x_{1}}{\beta_{2}+x_{1}}&-\frac{\beta_{1}x_{1}}{(\beta_{2}+x_{1})^{2}}\\ \vdots&\vdots\\ \frac{x_{n}}{\beta_{2}+x_{n}}&-\frac{\beta_{1}x_{n}}{(\beta_{2}+x_{n})^{2}}\end{bmatrix}.$ (29)
@@
$\widehat{\bm{\beta}}_{\alpha}-\bm{\beta}\xrightarrow{\mathcal{D}}\mathcal{N}_{2}\left(\bm{0},v_{1\alpha}\sigma^{2}\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)^{-1}\right),$
@@
$\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}=\begin{bmatrix}\sum_{i=1}^{n}a_{i}^{2}&-\sum_{i=1}^{n}a_{i}b_{i}\\ \\ -\sum_{i=1}^{n}a_{i}b_{i}&\sum_{i=1}^{n}b_{i}^{2}\end{bmatrix},$ (30)
with $a_{i}=\frac{x_{i}}{\beta_{2}+x_{i}}$, $b_{i}=\frac{\beta_{1}x_{i}}{(\beta_{2}+x_{i})^{2}}$. Upon simplifications, we get
$\widehat{\beta_{1}}_{\alpha}-\beta_{1}\xrightarrow{\mathcal{D}}\mathcal{N}\left(0,\sigma_{\beta_{1}}^{2}\right)\mbox{ and }\widehat{\beta_{2}}_{\alpha}-\beta_{2}\xrightarrow{\mathcal{D}}\mathcal{N}\left(0,\sigma_{\beta_{2}}^{2}\right),$
where $\sigma_{\beta_{1}}^{2}=\frac{v_{1\alpha}\sigma^{2}\sum_{i=1}^{n}b_{i}^{2}}{\sum_{i=1}^{n}a_{i}^{2}\sum_{i=1}^{n}b_{i}^{2}-\left(\sum_{i=1}^{n}a_{i}b_{i}\right)^{2}}$ and $\sigma_{\beta_{2}}^{2}=\frac{v_{1\alpha}\sigma^{2}\sum_{i=1}^{n}a_{i}^{2}}{\sum_{i=1}^{n}a_{i}^{2}\sum_{i=1}^{n}b_{i}^{2}-\left(\sum_{i=1}^{n}a_{i}b_{i}\right)^{2}}$.
In order to help verification of the required Conditions (R1)-(R2) for a given MM model, we have proved the following propositions; see Appendices B.2 and B.3 for their proofs.
Proposition 1.
Under the MM model given in (28), if the parameter space for $\beta_{2}$ is $[0,\infty)$ and $x_{i}>0$ for all $i$, then Condition (R1) is satisfied.
Proposition 2.
In order to satisfy Condition (14) in (R2) for the MM model given by (28), it is necessary that the values $x_{i}$’s are not all same, $n\geq 2$ and $\beta_{1}\neq 0$.
Note that, Proposition 1 provides a set of easily verifiable sufficient conditions for (R1) which are relatively stricter and not necessary for (R1). It may be possible to verify (R1) directly even if these sufficient conditions are not satisfied (although they are assumed to be true in most applications of the MM model). On the other hand, Proposition 2 provides a set of conditions which are necessary for (R2) to hold.
4.2 Influence function
@@
$IF_{i_{0}}(t_{i_{0}},T_{\alpha}^{\beta_{1}},\bm{F_{\theta}})=\\ \frac{(1+\alpha)^{\frac{3}{2}}(t_{i_{0}}-\mu(x_{i_{0}},\bm{\beta}))e^{-\frac{\alpha}{2\sigma^{2}}(t_{i_{0}}-\mu(x_{i_{0}},\bm{\beta}))^{2}}}{\det\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)}\frac{\beta_{1}^{2}x_{i_{0}}}{(\beta_{2}+x_{i_{0}})^{2}}\sum_{j=1}^{n}\frac{x_{j}^{2}(x_{i_{0}}-x_{j})}{(\beta_{2}+x_{j})^{4}},$ (31)
$IF_{i_{0}}(t_{i_{0}},T_{\alpha}^{\beta_{2}},\bm{F_{\theta}})=\\ \frac{(1+\alpha)^{\frac{3}{2}}(t_{i_{0}}-\mu(x_{i_{0}},\bm{\beta}))e^{-\frac{\alpha}{2\sigma^{2}}(t_{i_{0}}-\mu(x_{i_{0}},\bm{\beta}))^{2}}}{\det\left(\bm{\dot{\mu}(\beta)}^{T}\bm{\dot{\mu}(\beta)}\right)}\frac{\beta_{1}x_{i_{0}}}{(\beta_{2}+x_{i_{0}})^{2}}\sum_{j=1}^{n}\frac{x_{j}^{2}(x_{i_{0}}-x_{j})}{(\beta_{2}+x_{j})^{3}},$ (32)
when there is contamination only in $i_{0}$-th observation at the contamination point $t_{i_{0}}$. For infinitesimal contamination in all observations at the contamination points in $\bm{t}=(t_{1},\cdots,t_{n})^{T}$, these IFs are given by
$IF(\bm{t},T_{\alpha}^{\beta_{k}},\bm{F_{\theta}})=\sum_{i=1}^{n}IF_{i}(t_{i},T_{\alpha}^{\beta_{k}},\bm{F_{\theta}}),\ \ k=1,2.$ (33)
For illustrations, we have plotted these IFs of the MDPDEs under the MM-model (28), by generating data from the following two particular setups:
@@
  • Setup 2: $n=50,\ (\beta_{1},\beta_{2},\sigma)=(50,2,2),$ and $(x_{1},x_{2},\ldots,x_{n})$ is a random sample of size $n$ from Uniform$(0,40)$.
The resulting plots of the IFs for contaminations in all observations are presented in Figure 1 for both Setups 1-2, respectively; those obtained for contamination in particular observation are provided in Appendix C.1. To plot the IFs in (33) (in Figure 1) on a 2-D surface, we have considered same contamination point for all observations, i.e., the plots are obtained for the IFs at contamination points in $\bm{t}^{*}=(t,\ldots,t)$. From all these figures, it is clear that the IFs of the MDPDEs under the MM model are bounded for $\alpha>0$ and are redecending in nature for increasing values of $\alpha$. For $\alpha=0$, the IFs are unbounded as expected, since the MDPDE with $\alpha=0$ is the same as the non-robust MLE also for the MM models.
(a) $IF(\bm{t}^{*},T_{\alpha}^{\beta_{1}},\bm{F_{\theta}})$, under Setup 1
(b) $IF(\bm{t}^{*},T_{\alpha}^{\beta_{2}},\bm{F_{\theta}})$, under Setup 1
(c) $IF(\bm{t}^{*},T_{\alpha}^{\beta_{1}},\bm{F_{\theta}})$, under Setup 2
(d) $IF(\bm{t}^{*},T_{\alpha}^{\beta_{2}},\bm{F_{\theta}})$, under Setup 2
Figure 1: IFs of the MDPDEs under the MM model, with contamination in all observations, at the same contamination point $t$ [i.e., $\bm{t}^{*}=(t,\ldots,t)$]
@@ 4.3 Finite sample performances of the MDPDE
We now present an extensive simulation study to justify the improved performances of the MDPDEs of parameters of the MM model (28). For this purpose we have simulated data following the two setups (Setup 1 and Setup 2) as mentioned in the previous Section 4.2. Additionally, we have incorporated data contaminations by multiplying the observed values of response or covariate (or both) by a fixed number, say $C$, in a certain proportion (say $e_{c}$) of cases selected randomly from the full sample. We report the results for the choice of $C=5$ under Setup 1 and for $C=2$ under Setup 2.
For each setup, we present the empirical absolute bias (EBias) and the empirical mean square error (EMSE) of the MDPDE and different other robust estimators (as described in Appendix A.2) of the parameters of the MM model (28), based on $1000$ replications for pure data as well as contaminated data with different choices of the contamination proportion $e_{c}$. The results for $e_{c}=10\%$ (light contamination) and $e_{c}=40\%$ (heavy contamination) are reported in Tables 3-4, along with the same obtained for the pure data ($e_{c}=0\%$); the results obtained for $e_{c}=20\%$ and $30\%$ (moderate contamination) are reported in Appendix C.2 for brevity. In particular, we have compared the performances of the MDPDEs at different values of $\alpha$ with the OLS (which is the same as the MDPDE at $\alpha=0$), the MOM estimator, the KPS estimator with different choice of its tuning parameter $w$, and M-estimators based on the Huber’s and the Tukey’s loss/weight functions.
In addition, we have also computed the average prediction errors for the model fitted with different estimators which we define based only on the clean observations as given by
@@
[…]
$\left.r\mbox{-th replication,}\ i=1,2,\ldots,n\right\}$ and $\widehat{\bm{\beta}}^{(r)}$ be parameter estimate obtained at $r$-th replication.
Based on all these numerical results, we can easily see that the MDPDEs with suitable values of $\alpha$ perform the best among all the competitors considering all the cases of simulation setups. For lower contamination proportions, a close competitor of our MDPDE is the KPS estimator and the Tukey’s M estimator, but their performance deteriorates as contamination proportion increases. This deterioration is more significant when there are contamination in both response and covariate. Further, from Tables 5 and 10 it is clear that, for increasing proportion of contamination, the average prediction error gets minimized for a larger value of the MDPDE tuning parameter $\alpha$.
Outlier proportion 10% 40% 10% 40%
Empirical Bias
OLS 0.0110 0.0677 2.6532 2.3313 9.1277 2.1085 69.35 377.97 36.35 119.67
Empirical MSE
OLS 0.0422 0.2796 15.91 143.76 130.23 221.72 5931 180529 1358 15184
@@
MDPDE($\alpha$=0.05) 0.0056 0.0023 0.0521 0.0155 0.0697 0.6136 2.5806 2.1279 19.16 0.4343 0.0112 0.9795 24.08 4.5723 19.37
MDPDE($\alpha$=0.1) 0.0059 0.0025 0.0515 0.0168 0.0089 0.0212 0.4501 0.4815 4.0300 0.0430 0.0218 0.0012 5.8616 0.7556 6.4756
MDPDE($\alpha$=0.2) 0.0074 0.0030 0.0534 0.0007 0.0017 0.0272 0.0469 0.0288 0.1562 0.0167 0.0095 0.0192 0.1414 0.0778 0.2317
MDPDE($\alpha$=0.3) 0.0094 0.0036 0.0583 0.0047 0.0002 0.0205 0.0240 0.0144 0.2090 0.0084 0.0059 0.0158 0.1110 0.0537 0.2486
MDPDE($\alpha$=0.5) 0.0141 0.0048 0.0730 0.0089 0.0012 0.0049 0.0044 0.0063 0.4105 0.0004 0.0028 0.0026 0.0706 0.0355 0.4303
MDPDE($\alpha$=0.7) 0.0185 0.0058 0.0914 0.0126 0.0023 0.0097 0.0004 0.0044 0.6267 0.0056 0.0007 0.0107 0.0565 0.0292 0.6363
MDPDE($\alpha$=1) 0.0269 0.0081 0.1197 0.0206 0.0051 0.0257 0.0035 0.0031 0.9167 0.0159 0.0030 0.0257 0.0454 0.0242 0.9152
@@
MDPDE($\alpha$=0.05) 0.2963 0.0245 0.0445 0.5644 0.06908 4.9107 416.30 42.78 611.86 2.8726 0.0941 7.3089 1537.79 127.73 570.47
MDPDE($\alpha$=0.1) 0.2986 0.0248 0.0450 0.3367 0.02840 0.0548 56.44 4.6886 124.26 0.3459 0.0304 0.0603 488.85 85.12 184.64
MDPDE($\alpha$=0.2) 0.3071 0.0257 0.0474 0.3401 0.02841 0.0517 0.6918 0.0553 0.4673 0.3468 0.0293 0.0529 2.0120 0.1241 0.4688
MDPDE($\alpha$=0.3) 0.3203 0.0270 0.0512 0.3530 0.02962 0.0550 0.5915 0.0506 0.1591 0.3574 0.0302 0.0555 0.6441 0.0628 0.1901
MDPDE($\alpha$=0.5) 0.3561 0.0304 0.0615 0.3873 0.03285 0.0645 0.5914 0.0487 0.3042 0.3891 0.0332 0.0649 0.6338 0.0577 0.3275
MDPDE($\alpha$=0.7) 0.4013 0.0345 0.0738 0.4273 0.03652 0.0746 0.6063 0.0499 0.5570 0.4279 0.0368 0.0748 0.6431 0.0574 0.5724
MDPDE($\alpha$=1) 0.4777 0.0411 0.0920 0.4927 0.04275 0.0881 0.6240 0.0513 1.0455 0.4916 0.0429 0.0881 0.6542 0.0571 1.0430
@@
MDPDE($\alpha$=0.05) 0.961 3.147 54.59 6.333 18.58 3.829 16.13 315.87 21.39 242.94
MDPDE($\alpha$=0.1) 0.961 0.965 52.27 1.819 16.72 3.830 3.817 366.36 4.638 293.84
MDPDE($\alpha$=0.2) 0.962 0.960 47.68 0.957 13.97 3.835 3.819 146.64 3.819 221.57
MDPDE($\alpha$=0.3) 0.964 0.968 42.11 0.959 11.58 3.842 3.926 20.366 3.825 97.43
MDPDE($\alpha$=0.5) 0.971 0.965 4.665 0.964 4.348 3.861 3.844 3.765 3.842 4.427
MDPDE($\alpha$=0.7) 0.984 1.000 0.943 0.972 1.056 3.886 3.866 3.767 3.864 3.786
MDPDE($\alpha$=1) 0.995 0.987 0.963 0.985 0.943 3.929 3.902 3.776 3.900 3.792
We would like to particularly mention that the median of means (MOM) method fails to perform well for almost all the contamination schemes. So, we have not considered this MOM estimator for further real data applications presented in this paper.

### v2

@@ 4.1 Asymptotic properties
In order to find the asymptotic distributions of the MDPDEs of the parameters of the MM model, we follow the notation of Section 2.2 to get $\bm{\mu}(\bm{\beta})=\left(\frac{\beta_{1}x_{1}}{\beta_{2}+x_{1}},\ldots,\frac{\beta_{1}x_{n}}{\beta_{2}+x_{n}}\right)^{T}$ and hence
$\bm{\dot{\mu}}(\bm{\beta})=\begin{bmatrix}\frac{x_{1}}{\beta_{2}+x_{1}}&-\frac{\beta_{1}x_{1}}{(\beta_{2}+x_{1})^{2}}\\ \vdots&\vdots\\ \frac{x_{n}}{\beta_{2}+x_{n}}&-\frac{\beta_{1}x_{n}}{(\beta_{2}+x_{n})^{2}}\end{bmatrix}$. (29)
@@
$\left(\bm{\dot{\mu}}(\bm{\beta})^{T}\bm{\dot{\mu}}(\bm{\beta})\right)^{1/2}\left(\widehat{\bm{\beta}}_{\alpha}-\bm{\beta}\right)\xrightarrow{\mathcal{D}}\mathcal{N}_{2}\left(\bm{0},v_{1\alpha}\sigma^{2}\mathbb{I}_{2}\right),$
@@
$\bm{\dot{\mu}}(\bm{\beta})^{T}\bm{\dot{\mu}}(\bm{\beta})=\begin{bmatrix}\sum_{i=1}^{n}a_{i}^{2}&-\sum_{i=1}^{n}a_{i}b_{i}\\ -\sum_{i=1}^{n}a_{i}b_{i}&\sum_{i=1}^{n}b_{i}^{2}\end{bmatrix}$, (30)
with $a_{i}=\frac{x_{i}}{\beta_{2}+x_{i}}$, $b_{i}=\frac{\beta_{1}x_{i}}{(\beta_{2}+x_{i})^{2}}$. Upon simplification, we get
$\sqrt{n}\left(\widehat{\beta}_{1\alpha}-\beta_{1}\right)\xrightarrow{\mathcal{D}}\mathcal{N}\left(0,\sigma_{\beta_{1}}^{2}\right)\mbox{ and }\sqrt{n}\left(\widehat{\beta}_{2\alpha}-\beta_{2}\right)\xrightarrow{\mathcal{D}}\mathcal{N}\left(0,\sigma_{\beta_{2}}^{2}\right),$
where $\sigma_{\beta_{1}}^{2}=\lim\limits_{n\to\infty}\frac{nv_{1\alpha}\sigma^{2}\sum_{i=1}^{n}b_{i}^{2}}{\sum_{i=1}^{n}a_{i}^{2}\sum_{i=1}^{n}b_{i}^{2}-\left(\sum_{i=1}^{n}a_{i}b_{i}\right)^{2}}$ and $\sigma_{\beta_{2}}^{2}=\lim\limits_{n\to\infty}\frac{nv_{1\alpha}\sigma^{2}\sum_{i=1}^{n}a_{i}^{2}}{\sum_{i=1}^{n}a_{i}^{2}\sum_{i=1}^{n}b_{i}^{2}-\left(\sum_{i=1}^{n}a_{i}b_{i}\right)^{2}}$, but they are not independent in general.
In order to help verification of the required Conditions (R1)-(R2) for a given MM model, we have proved the following propositions; see Web Appendix B for their proofs.
Proposition 4.1.
Under the MM model given in (28), if the parameter space for $\beta_{2}$ is $0,\infty)$ and $x_{i}>0$ for all $i$, then Condition (R1) is satisfied.
Proposition 4.2.
In order to satisfy Condition (13) in (R2) for the MM model given by (28), it is necessary that the values $x_{i}$’s are not all same, $n\geq 2$, and $\beta_{1}\neq 0$.
Note that, Proposition 4.1 provides a set of easily verifiable sufficient conditions for (R1) which are relatively stricter and not necessary for (R1). It may be possible to verify (R1) directly even if these sufficient conditions are not satisfied (although they are assumed to be true in most applications of the MM model). On the other hand, Proposition 4.2 provides a set of conditions that are necessary for (R2) to hold.
4.2 Influence functions
@@
$IF_{i_{0}}(t_{i_{0}},T_{\alpha}^{\beta_{1}},\bm{F_{\theta}})=\frac{(1+\alpha)^{\frac{3}{2}}(t_{i_{0}}-\mu(x_{i_{0}},\bm{\beta}))e^{-\frac{\alpha}{2\sigma^{2}}(t_{i_{0}}-\mu(x_{i_{0}},\bm{\beta}))^{2}}}{\det\left(\bm{\dot{\mu}}(\bm{\beta})^{T}\bm{\dot{\mu}}(\bm{\beta})\right)}\frac{\beta_{1}^{2}x_{i_{0}}}{(\beta_{2}+x_{i_{0}})^{2}}\sum_{j=1}^{n}\frac{x_{j}^{2}(x_{i_{0}}-x_{j})}{(\beta_{2}+x_{j})^{4}},$ (31)
$IF_{i_{0}}(t_{i_{0}},T_{\alpha}^{\beta_{2}},\bm{F_{\theta}})=\frac{(1+\alpha)^{\frac{3}{2}}(t_{i_{0}}-\mu(x_{i_{0}},\bm{\beta}))e^{-\frac{\alpha}{2\sigma^{2}}(t_{i_{0}}-\mu(x_{i_{0}},\bm{\beta}))^{2}}}{\det\left(\bm{\dot{\mu}}(\bm{\beta})^{T}\bm{\dot{\mu}}(\bm{\beta})\right)}\frac{\beta_{1}x_{i_{0}}}{(\beta_{2}+x_{i_{0}})^{2}}\sum_{j=1}^{n}\frac{x_{j}^{2}(x_{i_{0}}-x_{j})}{(\beta_{2}+x_{j})^{3}},$ (32)
when there is contamination only in $i_{0}$-th observation at the contamination point $t_{i_{0}}$. Further, the corresponding IF of the minimum DPD functional $T_{\alpha}^{\sigma^{2}}$, denoted by $IF_{i_{0}}(t_{i_{0}},T_{\alpha}^{\sigma^{2}},\bm{F_{\theta}})$, remains the same as given in the expression (17), with $\mu(\bm{x}_{i_{0}},\bm{\beta})=\beta_{1}x_{i_{0}}/(\beta_{2}+x_{i_{0}})$. For infinitesimal contamination in all observations at the contamination points in $\bm{t}=(t_{1},\cdots,t_{n})^{T}$, these IFs are given by
$IF(\bm{t},T_{\alpha}^{\beta_{k}},\bm{F_{\theta}})=\sum_{i=1}^{n}IF_{i}(t_{i},T_{\alpha}^{\beta_{k}},\bm{F_{\theta}}),\ \ k=1,2,{\color[rgb]{0,0,0}\definecolor[named]{pgfstrokecolor}{rgb}{0,0,0}\pgfsys@color@gray@stroke{0}\pgfsys@color@gray@fill{0}\mbox{ and }IF(\bm{t},T_{\alpha}^{\sigma^{2}},\bm{F_{\theta}})=\sum_{i=1}^{n}IF_{i}(t_{i},T_{\alpha}^{\sigma^{2}},\bm{F_{\theta}})}.$ (33)
(a) $IF(\bm{t}^{*},T_{\alpha}^{\beta_{1}},\bm{F_{\theta}})$, under Setup 1
(b) $IF(\bm{t}^{*},T_{\alpha}^{\beta_{1}},\bm{F_{\theta}})$, under Setup 2
(c) $IF(\bm{t}^{*},T_{\alpha}^{\beta_{2}},\bm{F_{\theta}})$, under Setup 1
(d) $IF(\bm{t}^{*},T_{\alpha}^{\beta_{2}},\bm{F_{\theta}})$, under Setup 2
(e) $IF(\bm{t}^{*},T_{\alpha}^{\sigma^{2}},\bm{F_{\theta}})$, under Setup 1
(f) $IF(\bm{t}^{*},T_{\alpha}^{\sigma^{2}},\bm{F_{\theta}})$, under Setup 2
Figure 1: IFs of the MDPDEs of the parameters $\beta_{1}$, $\beta_{2}$, and $\sigma^{2}$ in the MM model, with contamination in all observations, at the same contamination point, i.e., for $\bm{t}^{*}=(t,\ldots,t)^{T}$.
For illustrations, we have plotted these IFs of the MDPDEs under the MM model, by generating data from the following two particular setups:
@@
  • Setup 2: $n=50,\ (\beta_{1},\beta_{2},\sigma)=(50,2,2),$ and $(x_{1},x_{2},\ldots,x_{n})$ is a random sample of size $n$ from Uniform(0,40)(0,40).
The resulting IFs for contaminations in all observations are presented in the Figure 1 for both Setups 1-2, where the contamination points are taken as $\bm{t}^{*}=(t,\ldots,t)$; those obtained for contamination in particular observation are provided in Figures C.1-C.2 of Web Appendix C. From all these figures, it is clear that the IFs of the MDPDEs under the MM model are bounded for $\alpha>0$ and are redecending in nature for increasing values of $\alpha$. For $\alpha=0$, the IFs are unbounded as expected, since the MDPDE with $\alpha=0$ is the same as the non-robust MLE also for the MM models.
@@ 4.3 Finite sample performances of the MDPDE
We now present an extensive simulation study to justify the improved performances of the MDPDEs of parameters of the MM model (28). For this purpose, we have simulated data following the two setups (Setup 1 and Setup 2) as mentioned in the Section C.1. Note that the choice of true parameter values for Setup 1, i.e., $\bm{\beta}=(5,1)$, is made directly following Tabatabai et al. [39]. The parameter values of the second setup are chosen to be moderately large, namely (50, 2), to closely reproduce real-life examples following the discussions in Marasovic et al. [31]. Additionally, we have incorporated data contamination by multiplying the observed values of response or covariate (or both) by a fixed number, say $C$, in a certain proportion (say $e_{c}$) of cases selected randomly from the full sample. We report the results for the choice of $C=5$ under Setup 1 and for $C=2$ under Setup 2. In each case, we have used the R function optim [^2] https://stat.ethz.ch/R-manual/R-devel/library/stats/html/optim.html (with the BFGS method) to compute the MDPDEs by minimizing the objective function in (5), where the initial values of the parameters are chosen to be their true values; this helped us to compare the actual behavior of the estimates, avoiding any numerical instability and the effects of initial values.
For each setup, we present the absolute empirical bias (EBias) and the empirical mean square error (EMSE) of the MDPDE and different other robust estimators of the parameters of the MM model (28), based on $1000$ replications. For a generic model parameter $\theta$ (which is either of $\beta_{1}$, $\beta_{2}$ or $\sigma$ in our case) having true value $\theta_{0}$, the summary measures EBias and EMSE are computed as
$EBias=\left|\frac{1}{1000}\sum_{r=1}^{1000}\widehat{\theta}^{(r)}-\theta_{0}\right|,\mbox{ and }EMSE=\frac{1}{1000}\sum_{r=1}^{1000}\left(\widehat{\theta}^{(r)}-\theta_{0}\right)^{2},$
where $\widehat{\theta}^{(1)},\ldots,\widehat{\theta}^{(1000)}$ denote the estimated values of the parameter $\theta$ across 1000 replications. We have repeated this exercise for pure data under both the setups, as well as contaminated data with different choices of the contamination proportion $e_{c}$. The results for $e_{c}=0\%$ (pure data), $e_{c}=10\%$ (light contamination) and $e_{c}=40\%$ (heavy contamination) are reported in Table 1 (for Setup 2) and Table C.4 of Web Appendix C (for Setup 1); the results obtained for $e_{c}=20\%$ and $30\%$ (moderate contamination) are reported in Tables C.5-C.6 of Web Appendix C for both setups.
In all these simulations, we have compared the performances of the MDPDEs at different values of $\alpha$ with the OLS (which is the same as the MDPDE at $\alpha=0$), the MOM estimator, the KPS estimator with various choices of its tuning parameter $w$, and M-estimators based on the Huber’s and the Tukey’s loss/weight functions (described in Web Appendix A.2). These competing estimators are computed following the existing algorithms provided in their respective reference papers, and a comparison of the runtime complexity of these estimates with our proposed MDPDEs is provided in Web Appendix C.3 (for the simulation Setup 2).
In addition, we have also computed the average prediction errors for the model fitted with different estimators, which we define based only on the clean observations as given by
@@
where $\mathcal{A}_{r}=\{i:$ $i$-th data-point is not contaminated in the dataset generated at $r$-th replication, $i=1,2,\ldots,n\}$ and $\widehat{\bm{\beta}}^{(r)}$ be the parameter estimate obtained at $r$-th replication for a particular method. This measures the expected prediction accuracy obtained from a single dataset using any of the robust estimators.
Table 1: absolute empirical bias and empirical MSE obtained by different estimators in simulations with Setup 2 and different amounts of contamination (the minimum values obtained for each parameter are highlighted in bold font)
@@
MDPDE($\alpha$=0.05) 0.0056 0.0023 0.0521 0.0139 0.0684 0.5923 14.09 0.0759 13.80 0.4244 0.0106 0.9497 41.15 9.9069 13.49
MDPDE($\alpha$=0.1) 0.0060 0.0025 0.0515 0.0168 0.0089 0.0212 2.4492 0.0940 2.9793 0.0429 0.0218 0.0012 11.45 2.5475 4.4030
MDPDE($\alpha$=0.2) 0.0075 0.0030 0.0534 0.0001 0.0019 0.0268 0.0468 0.0288 0.1563 0.0166 0.0094 0.0192 0.1417 0.0779 0.2318
MDPDE($\alpha$=0.3) 0.0094 0.0036 0.0583 0.0048 0.0001 0.0206 0.0240 0.0144 0.2089 0.0084 0.0059 0.0158 0.1111 0.0538 0.2487
MDPDE($\alpha$=0.5) 0.0144 0.0048 0.0733 0.0089 0.0012 0.0049 0.0043 0.0063 0.4106 0.0003 0.0028 0.0026 0.0706 0.0356 0.4304
MDPDE($\alpha$=0.7) 0.0186 0.0059 0.0914 0.0127 0.0023 0.0097 0.0005 0.0043 0.6268 0.0057 0.0007 0.0107 0.0566 0.0292 0.6364
MDPDE($\alpha$=1) 0.0270 0.0081 0.1198 0.0207 0.0051 0.0257 0.0038 0.0030 0.9168 0.0160 0.0030 0.0257 0.0453 0.0241 0.9153
@@
MDPDE($\alpha$=0.05) 0.2964 0.0245 0.0445 0.5847 0.0678 4.7267 294.64 0.8267 268.86 2.8974 0.0959 7.0403 2148.77 153.35 212.34
MDPDE($\alpha$=0.1) 0.2986 0.0248 0.0450 0.3369 0.0284 0.0548 45.49 0.1394 56.41 0.3460 0.0304 0.0603 561.17 36.60 69.36
MDPDE($\alpha$=0.2) 0.3073 0.0257 0.0474 0.3402 0.0284 0.0516 0.6943 0.0554 0.4675 0.3471 0.0293 0.0529 2.0103 0.1239 0.4676
MDPDE($\alpha$=0.3) 0.3206 0.0270 0.0512 0.3535 0.0297 0.0550 0.5938 0.0508 0.1592 0.3579 0.0302 0.0555 0.6463 0.0630 0.1902
MDPDE($\alpha$=0.5) 0.3572 0.0304 0.0615 0.3883 0.0329 0.0645 0.5941 0.0489 0.3043 0.3902 0.0333 0.0649 0.6364 0.0579 0.3276
MDPDE($\alpha$=0.7) 0.4025 0.0346 0.0738 0.4286 0.0366 0.0746 0.6092 0.0501 0.5571 0.4292 0.0369 0.0748 0.6462 0.0576 0.5725
MDPDE($\alpha$=1) 0.4791 0.0412 0.0920 0.4945 0.0429 0.0882 0.6272 0.0515 1.0455 0.4935 0.0430 0.0881 0.6574 0.0573 1.0431
@@
MDPDE($\alpha$=0.05) 0.961 3.147 54.59 6.333 18.58 3.829 4.414 202.00 4.785 211.52
MDPDE($\alpha$=0.1) 0.961 0.965 52.27 1.819 16.72 3.830 3.817 40.69 3.826 65.29
MDPDE($\alpha$=0.2) 0.962 0.960 47.68 0.957 13.97 3.835 3.820 3.924 3.823 4.056
MDPDE($\alpha$=0.3) 0.964 0.968 42.11 0.959 11.58 3.842 3.826 3.763 3.828 3.802
MDPDE($\alpha$=0.5) 0.971 0.965 4.665 0.964 4.348 3.861 3.844 3.760 3.846 3.789
MDPDE($\alpha$=0.7) 0.984 1.000 0.943 0.972 1.056 3.886 3.866 3.768 3.867 3.789
MDPDE($\alpha$=1) 0.995 0.987 0.963 0.985 0.943 3.929 3.902 3.776 3.903 3.795
Based on all these numerical results, we can easily see that the MDPDEs with suitable values of $\alpha$ perform the best among all the competitors considering all the cases of simulation setups. For lower contamination proportions, a close competitor of our MDPDE is the KPS estimator and the Tukey’s M estimator, but their performance deteriorates as contamination proportion increases. This deterioration is more significant when there are contamination in both response and covariate. Further, for increasing proportion of contamination, the average prediction error gets minimized for a larger value of the tuning parameter $\alpha$ (refer to Table 2 and Table C.3 in Web Appendix C).
We would like to particularly mention that the median of means (MOM) method fails to perform well for almost all the contamination schemes. So, we have not considered the MOM estimator for further real data applications presented in this paper.